# NuSegDG: Integration of Heterogeneous Space and Gaussian Kernel for Domain-Generalized Nuclei Segmentation

### v2

@@ 4.1.2 Implementation Details
We conduct our experiments on two parallel NVIDIA Tesla P40 GPUs (48GB), utilizing PyTorch 1.13.0, Python 3.10, and CUDA 11.7. We maintain consistent training settings and configurations across all experiments to ensure fairness and reproducibility. For the optimizer, we employ Adam with a batch size of 2 and train models for 100 epochs. The initial learning rate is set to 0.0001 and is adjusted using an exponential decay strategy with a decay factor of 0.98. The loss coefficient $\alpha$ and $\beta$ are set to 0.8 and 0.2 during the training. In our proposed NuSegDG framework, the number of heterogeneous space $N$ is set to 2. All images are resized to $1024\times 1024$. To save computational costs, the ViT-B [10] is considered as the image encoder for all SAM-based frameworks. For the TSM-Decoder, we select the pre-trained StarDIST [55] as our auxiliary neural network to facilitate accurate instance segmentation without manual morphological shape refinement. We utilize the single-point prompt to fine-tune all SAM-based architectures. The point is generated using the connectedComponents in OpenCV, which is the centroid of each nucleus instance.
@@ 4.2 Evaluation Metrics
In our experiments, we first evaluate the performance of models on the semantic segmentation task using four common metrics: Dice coefficient, mean Intersection over Union (mIoU), F1-score, and Hausdorff Distance (HD). Then, we adopt four extra metrics: Aggregated Jaccard Index (AJI), Detection Quality (DQ), Segmentation Quality (SQ), and Panoptic Quality (PQ), to make comparisons on the instance segmentation task. Especially, the best and second-best performance values are highlighted in bold and underlined. For each task, we use two different evaluation protocols: domain generalization and adaptability evaluation.
4.2.1 Domain Generalization Evaluation
We employ a standard leave-one-domain-out strategy [56] to conduct the domain generalization evaluation. Specifically, the model is trained on a training set $\mathcal{S}$ of $K-1$ source domains, where each source domain represents a different data distribution, and then evaluated on the remaining unseen target domains $\mathcal{T}$, e.g, $\mathcal{S}=\{\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3}\}$, $\mathcal{T}=\mathcal{S}_{4}$.
4.2.2 Adaptability Evaluation
@@
In this protocol, we perform a fully supervised learning where all four datasets (i.e., $\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3},\mathcal{S}_{4}$) are considered as seen domains. We randomly divide all datasets into three sets: training, validation, and testing, in the conventional ratio of 8:1:1. The model is evaluated on the testing set of each dataset individually. The reason for conducting this protocol is to assess the adaptability of each model across different domains. Moreover, we display the performance gap between our domain generalization approach and traditional fully supervised methods. This comparison further demonstrates the effectiveness of NuSegDG on domain generalization.
Table 3: Domain-generalized Comparison with state-of-the-arts on nuclei instance segmentation.
Datasets $\mathcal{T}=\mathcal{S}_{1}$ $\mathcal{T}=\mathcal{S}_{2}$ $\mathcal{T}=\mathcal{S}_{3}$ $\mathcal{T}=\mathcal{S}_{4}$
Methods AJI DQ SQ PQ AJI DQ SQ PQ AJI DQ SQ PQ AJI DQ SQ PQ
U-Net [20] 6.35 4.94 10.74 3.41 31.13 28.70 66.70 19.86 14.24 16.19 48.22 10.88 42.35 47.64 72.10 34.51
Mask-RCNN [57] 7.73 7.26 12.82 5.08 30.06 30.71 66.49 21.74 11.30 14.05 48.53 9.37 43.57 48.98 71.73 35.30
StarDIST [55] 8.17 6.55 11.90 4.52 35.58 39.40 70.08 28.49 18.66 24.19 56.50 16.53 43.81 49.61 71.79 35.71
Hover-Net [7] 5.42 2.52 16.37 1.75 42.07 53.42 72.96 39.43 23.11 25.50 65.28 17.73 41.10 40.63 64.17 30.48
TSFD-Net [34] 6.03 5.23 11.53 3.68 41.37 49.92 71.41 35.87 12.48 17.02 51.67 11.36 37.95 45.16 71.56 32.50
CellPose [31] 15.75 19.00 36.72 13.98 20.56 27.47 73.04 20.18 38.70 54.55 71.16 42.53 41.84 54.39 72.58 41.27
CPP-Net [8] 8.63 6.58 12.04 4.53 41.07 47.81 72.39 34.93 17.42 21.94 60.86 14.66 41.09 47.78 72.33 34.72
CellViT [9] 4.39 4.95 12.00 3.66 48.44 63.31 73.85 49.17 47.30 64.00 64.76 45.53 43.94 57.23 67.57 42.36
SAM [10] 41.95 39.24 63.75 30.00 47.42 55.73 71.85 40.41 53.92 57.56 72.24 43.79 41.00 48.43 73.03 35.62
Med-SA [43] 57.64 55.06 74.90 44.42 50.59 62.06 74.10 46.23 56.60 63.26 72.98 47.94 51.40 61.72 73.50 45.11
SAMed [47] 51.15 46.98 69.71 36.92 49.47 61.00 73.58 45.10 54.59 59.69 73.20 44.14 49.34 59.31 73.33 43.65
SAMUS [44] 60.98 61.41 77.80 50.17 50.62 62.17 74.11 46.32 57.27 67.65 72.60 50.32 51.34 62.07 73.40 45.68
SAM-CL [41] 52.76 51.63 71.08 41.17 47.84 57.26 72.99 42.00 55.29 60.79 73.49 46.30 49.36 59.26 73.59 43.78
PromptNucSeg [13] 50.89 48.66 70.44 38.87 48.82 59.06 72.24 43.03 54.14 57.31 72.79 43.78 45.49 44.44 71.73 32.10
Table 4: Adaptability Comparison with state-of-the-arts on nuclei instance segmentation.
Datasets $\mathcal{S}_{1}$ $\mathcal{S}_{2}$ $\mathcal{S}_{3}$ $\mathcal{S}_{4}$
Methods AJI DQ SQ PQ AJI DQ SQ PQ AJI DQ SQ PQ AJI DQ SQ PQ
U-Net [20] 63.49 74.81 81.72 61.45 50.27 61.75 74.74 46.23 51.34 60.73 75.32 45.88 46.43 55.90 75.80 42.38
Mask-RCNN [57] 63.32 75.03 81.05 61.44 45.32 55.98 74.16 41.61 43.84 54.79 74.71 41.35 46.72 56.16 76.25 42.87
StarDIST [3] 63.38 74.78 80.51 60.90 54.95 68.24 74.36 50.87 45.26 56.15 76.11 42.76 46.57 53.82 74.84 40.29
Hover-Net [7] 61.59 61.04 79.48 50.03 54.97 71.29 75.68 54.02 25.10 25.14 69.16 17.75 36.35 34.98 71.59 25.21
TSFD-Net [34] 62.25 72.67 80.14 59.06 54.16 67.98 74.55 50.78 42.93 53.77 75.97 40.85 47.64 58.15 75.20 43.79
CellPose [31] 66.77 80.21 82.54 66.93 20.62 27.82 74.08 20.72 45.21 62.82 76.88 48.40 36.46 45.90 74.92 34.61
CPP-Net [8] 63.60 74.72 81.51 61.48 52.29 66.21 73.94 49.03 56.00 68.20 77.75 53.05 47.36 56.03 76.00 43.21
CellViT [9] 60.51 73.92 84.10 63.25 57.91 77.35 77.19 60.54 53.16 68.72 77.65 54.83 40.87 58.13 76.16 44.38
SAM [10] 73.32 80.21 83.08 67.16 55.51 69.70 75.24 52.51 55.16 54.44 77.74 42.45 41.66 50.71 75.71 38.77
Med-SA [43] 74.17 84.89 81.95 70.03 64.53 78.64 75.45 59.42 64.59 67.92 77.11 52.89 50.35 59.52 73.80 44.26
SAMed [47] 71.95 82.30 80.04 66.42 62.62 75.52 75.70 57.29 63.30 65.15 75.09 49.09 49.14 56.94 73.94 42.42
SAMUS [44] 73.70 86.97 80.88 70.72 67.98 83.09 76.87 63.93 63.34 76.56 77.23 60.84 51.25 63.53 75.79 48.35
SAM-CL [41] 73.20 86.40 80.89 70.37 63.60 78.49 74.84 58.84 60.11 67.01 77.19 51.90 50.42 61.28 74.58 45.83
PromptNucSeg [13] 74.26 85.13 82.25 70.31 64.53 79.03 76.32 60.41 61.94 72.44 77.46 56.98 52.31 62.11 73.23 45.71
@@ 4.3 Comparison on Nuclei Semantic Segmentation
To comprehensively assess our NuSegDG, we first compare it with state-of-the-art frameworks on nuclei semantic segmentation. As illustrated in Table 1, in the domain generalization evaluation, TransUNet [28] achieves leading results among previous U-shape segmentation algorithms due to its large model capacity. Benefiting from pre-training on the large-scale dataset, PEFT SAMs [43, 47, 44, 41, 52, 53, 54] display better performance than these task-specific models. In contrast, our NuSegDG surpasses the second-best SAMUS by a significant mIoU increase of 2.62%, 3.07%, 2.55%, and 4.51% on these four target domains, respectively. Compared to the prompt-free SAMs, NuSegDG presents a mIoU rise of 5.01% to 12.57%.
Moreover, Table 2 provides results of the adaptability evaluation. We observe that previous U-shape architectures show remarkable performance gains in the seen domain but are inferior to PEFT SAMs. Our NuSegDG achieves superior performance on these four datasets, with the best mIoU of 87.46%, 76.06%, 78.93% and 73.44%, respectively. On the other hand, the domain-generalized NuSegDG in Table 1 demonstrates competitive performance on $\mathcal{S}_{2}$, $\mathcal{S}_{3}$ and $\mathcal{S}_{4}$ domains compared to fully-supervised U-shape and SAM-based architectures in Table 2. We provide the visualization results in Fig. 2. Consequently, these comparisons validate the superiority of our NuSegDG on domain-generalized nuclei semantic segmentation tasks and better adaptability in diverse nuclei domains.
Figure 2: Qualitative comparison with state-of-the-art task-specific models and medical SAMs on domain-generalized nuclei semantic segmentation across four domains: DSB-2018 [3], MoNuSeg-2018 [4], TNBC [5] and CryoNuSeg [6].
Figure 3: Qualitative comparison with state-of-the-art task-specific models and medical SAMs on domain-generalized nuclei instance segmentation across four domains: DSB-2018 [3], MoNuSeg-2018 [4], TNBC [5] and CryoNuSeg [6].
@@ 4.4 Comparison on Nuclei Instance Segmentation
To further evaluate our NuSegDG in nuclei instance segmentation tasks, we perform the comparison with advanced nuclei instance segmentation frameworks on four different nuclei domains. Firstly, Table 3 presents experimental results under the domain generalization evaluation. It is revealed that previous morphological post-processing algorithms [7, 34, 31, 8, 9] show poor generalization capabilities on the $\mathcal{S}_{1}$ domain. On the contrary, SAMUS [44] performs better than these methods by achieving a remarkable PQ of 50.17%, 46.32%, 50.32% and 45.68% on four domains, respectively. Our NuSegDG outperforms it with a significant PQ increase of 7.90%, 8.31%, 1.29%, and 2.83%, respectively.
@@
Furthermore, we provide the adaptability evaluation result in Table 4. It is demonstrated that PEFT SAMs outperform morphological post-processing algorithms in four nuclei datasets. For example, PromptNucSeg [13] has a 6.62% AJI increase over CellViT [9] on the $\mathcal{S}_{2}$ domain. In contrast, our NuSegDG framework achieves the best AJI of 77.91%, 69.81%, 73.08%, and 53.33%, respectively, on the four datasets, and performs better than the state-of-the-art methods in the other three evaluation metrics. The quantitative comparison is presented in Fig. 3. As a result, these results reveal a significant performance advantage of our NuSegDG over current medical foundation models and task-specific architectures on domain-generalized nuclei semantic and instance segmentation tasks.

### v3

@@ 4.1.2 Implementation Details
We conduct our experiments on two parallel NVIDIA Tesla P40 GPUs (48GB), utilizing PyTorch 1.13.0, Python 3.10, and CUDA 11.7. We maintain consistent training settings and configurations across all experiments to ensure fairness and reproducibility. For the optimizer, we employ Adam with a batch size of 2 and train models for 100 epochs. The initial learning rate is set to 0.0001 and is adjusted using an exponential decay strategy with a decay factor of 0.98. The loss coefficient $\alpha$ and $\beta$ are set to 0.8 and 0.2 during the training. In our proposed NuSegDG framework, the number of heterogeneous space $N$ is set to 2. All images are resized to $1024\times 1024$. To save computational costs, the ViT-B [kirillov2023segment] is considered as the image encoder for all SAM-based frameworks. For the TSM-Decoder, we select the pre-trained StarDIST [schmidt2018cell] as our auxiliary neural network to facilitate accurate instance segmentation without manual morphological shape refinement. We utilize the single-point prompt to fine-tune all SAM-based architectures. The point is generated using the connectedComponents in OpenCV, which is the centroid of each nucleus instance. For the fluorescence data (which are of single-channel), we replicate the single channel to create an RGB-like input by utilizing the cvtColor in OpenCV.
@@ 4.2 Evaluation Metrics
In our experiments, we first evaluate the performance of models on the semantic segmentation task using four common metrics: Dice coefficient, mean Intersection over Union (mIoU), F1-score, and Hausdorff Distance (HD). Then, we adopt four extra metrics: Aggregated Jaccard Index (AJI), Detection Quality (DQ), Segmentation Quality (SQ), and Panoptic Quality (PQ), to make comparisons on the instance segmentation task, defined by [horst2024cellvit]:
$PQ=\underbrace{\frac{|TP|}{|TP|+\frac{1}{2}|FP|+\frac{1}{2}|FN|}}_{\text{Detection Quality (DQ)}}\times\underbrace{\frac{\sum_{(G,P)\in TP}\text{IoU}(G,P)}{|TP|}}_{\text{Segmentation Quality (SQ)}},$ (13)
where $G$ is the ground truth and $P$ is the prediction segmentation mask. It is important to note that the IoU threshold plays a crucial role in matching predicted instances with ground truth instances, and in our experiments, we have set this threshold to the default value of 0.5 (which is the value that it strikes a balance between being lenient enough to capture true positives and strict enough to penalize poorly placed predictions). In addition, the best and second-best performance values are highlighted in bold and underlined. For each task, we use two different evaluation protocols: domain generalization and adaptability evaluation.
Figure 2: Qualitative comparison with state-of-the-art task-specific models and medical SAMs on generalized nuclei semantic segmentation across four target domains: DSB-2018 [caicedo2019nucleus], MoNuSeg-2018 [kumar2017dataset], TNBC [naylor2018segmentation] and CryoNuSeg [mahbod2021cryonuseg].
Figure 3: Qualitative comparison with state-of-the-art task-specific models and medical SAMs on generalized nuclei instance segmentation across four target domains: DSB-2018 [caicedo2019nucleus], MoNuSeg-2018 [kumar2017dataset], TNBC [naylor2018segmentation] and CryoNuSeg [mahbod2021cryonuseg].
4.2.1 Source Domain Generalization Evaluation
@@
In this protocol, we perform a fully supervised learning where all four datasets (i.e., $\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3},\mathcal{S}_{4}$) are considered as seen domains. We randomly divide all datasets into three sets: training, validation, and testing, in the conventional ratio of 8:1:1. The model is evaluated on the testing set of each dataset individually. The reason for conducting this protocol is to assess the adaptability of each model across different domains. Moreover, we display the performance gap between our domain generalization approach and traditional fully supervised methods. This comparison further demonstrates the effectiveness of NuSegDG on domain generalization tasks.
4.2.2 Target Domain Generalization Evaluation
We employ a standard leave-one-domain-out strategy [zhou2022domain] to conduct the domain generalization evaluation. Specifically, the model is trained on a training set $\mathcal{S}$ of $K-1$ source domains, where each source domain represents a different data distribution, and then evaluated on the remaining unseen target domains $\mathcal{T}$, e.g, $\mathcal{S}=\{\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3}\}$, $\mathcal{T}=\mathcal{S}_{4}$.
@@ 4.3 Comparison on Nuclei Semantic Segmentation
To comprehensively assess our NuSegDG, Table 1 provides generalization evaluation results on source domains. We observe that previous U-shape architectures show remarkable performance gains in the seen domain but are inferior to PEFT SAMs. Our NuSegDG achieves superior performance on these four datasets, with the best mIoU of 87.46%, 76.06%, 78.93% and 73.44%, respectively. On the other hand, the domain-generalized NuSegDG in Table 2 demonstrates competitive performance on $\mathcal{S}_{2}$, $\mathcal{S}_{3}$ and $\mathcal{S}_{4}$ domains compared to fully-supervised U-shape and SAM-based architectures in Table 1. We provide the visualization results in Fig. 2.
Moreover, we compare it with state-of-the-art frameworks on nuclei semantic segmentation. As illustrated in Table 2, in the target domain generalization evaluation, TransUNet [chen2024transunet] achieves leading results among previous U-shape segmentation algorithms due to its large model capacity. Benefiting from pre-training on the large-scale dataset, PEFT SAMs [wu2023medical, zhang2023customized, lin2023samus, zhong2024convolution, na2024segment, gu2024lesam, cheng2024unleashing] display better performance than these task-specific models. In contrast, our NuSegDG surpasses the second-best SAMUS by a significant mIoU increase of 2.62%, 3.07%, 2.55%, and 4.51% on these four target domains, respectively. Compared to the prompt-free SAMs, NuSegDG presents a mIoU rise of 5.01% to 12.57%. Consequently, these comparisons validate the superiority of our NuSegDG on domain-generalized nuclei semantic segmentation tasks in diverse nuclei domains.
Table 6: Comparison of inference time with the vanilla Point Prompt mode.
Datasets $\mathcal{S}_{1}$ $\mathcal{S}_{2}$ $\mathcal{S}_{3}$ $\mathcal{S}_{4}$
Nuclei (Avg.) 30 510 99 157
Vanilla Point Prompt 11.19s 61.49s 12.64s 19.92s
$+$ Auxiliary Neural Network 10.31s 11.62s 10.55s 10.86s
Table 7: Comparison with the original point prompt mode on nuclei semantic segmentation (Source Domain Generalization).
Datasets $\mathcal{S}_{1}$ $\mathcal{S}_{2}$ $\mathcal{S}_{3}$ $\mathcal{S}_{4}$
Prompt Types Dice mIoU F1 HD Dice mIoU F1 HD Dice mIoU F1 HD Dice mIoU F1 HD
1 point 92.09 85.62 92.20 58.29 84.45 73.17 84.44 59.96 85.53 74.76 85.48 85.28 83.08 71.32 83.43 90.49
3 points 92.66 86.78 92.71 60.49 84.38 73.12 84.41 59.64 86.77 76.67 86.62 101.42 82.85 71.04 82.98 91.97
5 points 92.93 87.08 92.95 41.74 84.40 73.11 84.43 39.01 86.99 77.03 86.92 51.73 84.09 72.86 84.18 86.73
7 points 93.08 87.26 93.11 53.27 85.06 74.07 85.09 55.81 87.57 77.94 87.50 59.45 84.22 72.93 84.30 83.39
1 point + GK 93.17 87.46 93.35 33.18 86.37 76.06 86.40 44.44 88.20 78.93 87.03 49.69 84.59 73.44 84.75 64.61
Table 8: Comparison of pretrained and fine-tuned auxiliary network on nuclei instance segmentation (Source Domain Generalization).
Datasets $\mathcal{S}_{1}$ $\mathcal{S}_{2}$ $\mathcal{S}_{3}$ $\mathcal{S}_{4}$
Methods Tuned AJI DQ SQ PQ AJI DQ SQ PQ AJI DQ SQ PQ AJI DQ SQ PQ
Auxiliary Network 77.91 88.88 85.47 76.31 69.81 88.66 77.68 68.88 73.08 85.33 78.15 66.84 53.33 63.64 76.87 49.11
78.50 89.20 86.00 77.10 70.40 89.50 78.90 69.80 73.80 85.90 78.70 67.50 54.00 64.20 77.50 50.00
@@ 4.4 Comparison on Nuclei Instance Segmentation
To further evaluate our NuSegDG in nuclei instance segmentation tasks, we provide the source domain generalization comparison result in Table 3. It is demonstrated that PEFT SAMs outperform morphological post-processing algorithms in four nuclei datasets. For example, PromptNucSeg [shui2023unleashing] has a 6.62% AJI increase over CellViT [horst2024cellvit] on the $\mathcal{S}_{2}$ domain. In contrast, our NuSegDG framework achieves the best AJI of 77.91%, 69.81%, 73.08%, and 53.33%, respectively, on the four datasets, and performs better than the state-of-the-art methods in the other three evaluation metrics.
Furthermore, we perform the comparison with advanced nuclei instance segmentation frameworks on four different nuclei domains. Firstly, Table 4 presents experimental results under the target domain generalization evaluation. It is revealed that previous morphological post-processing algorithms [graham2019hover, ilyas2022tsfd, stringer2021cellpose, chen2023cpp, horst2024cellvit] show poor generalization capabilities on the $\mathcal{S}_{1}$ domain. On the contrary, SAMUS [lin2023samus] performs better than these methods by achieving a remarkable PQ of 50.17%, 46.32%, 50.32% and 45.68% on four domains, respectively. Our NuSegDG outperforms it with a significant PQ increase of 7.90%, 8.31%, 1.29%, and 2.83%, respectively. The quantitative comparison is presented in Fig. 3. As a result, these results reveal a significant performance advantage of our NuSegDG over current medical foundation models and task-specific architectures on domain-generalized nuclei semantic and instance segmentation tasks.
4.5 Ablation Study
To investigate the effectiveness of the individual components within the NuSegDG framework, we conduct an ablation study on domain-generalized nuclei instance segmentation, as summarized in Table 5. This study sequentially enables or disables the HS-Adapter $M_{1}$, GKP-Encoder $M_{2}$, and TSM-Decoder $M_{3}$ to evaluate their impact on the performance of the average AJI, DQ, SQ, and PQ metrics. Firstly, we consider the standard fine-tuned SAM ($1^{st}$ row) as the ablation baseline. By respectively embedding the HS-Adapter ($2^{nd}$ row), GKP-Encoder ($3^{rd}$ row) and TSM-Decoder ($4^{th}$ row), the performance is raised with the average AJI of 5.17%, 2.79%, 2.07%, and the average PQ of 6.61%, 3.17%, 1.58%. When we combine HS-Adapter with GKP-Encoder ($5^{th}$ row), the performance of the model is further improved, with the average AJI of 56.29 and PQ of 50.38% on the four domains. This result proves that these two modules can promote the domain generalization capability in nuclei instance segmentation. By comparing $6^{th}$ and $7^{th}$ rows with $2^{nd}$ and $3^{rd}$ rows, the TSM-Decoder demonstrates significant performance gains while eliminating the demand for manual morphological refinement. Finally, our NuSegDG framework ($8^{th}$ row) integrates all three modules and achieves the best performance on all metrics, with an average AJI of 58.84%, an average DQ of 69.62%, an average SQ of 75.50%, and an average PQ of 53.21%. This full configuration significantly outperforms the others, emphasizing the synergistic benefits of incorporating all modules. This result highlights the importance of each component in enhancing the generalization capability of NuSegDG across different nuclei image domains.
Moreover, nuclei semantic and instance segmentation tasks meet two different requirements for clinical applications, such as disease area calculation and nuclei counting. Based on experimental results, we can observe that existing automatic SAMs show competitive generalization performance on source domains but are inferior to the SAMs with point prompts on target domains. On the other hand, although the vanilla point prompt performs better in domain generalization, it predicts cell instances one by one which is time-consuming for dense cell maps (e.g., whole slide imaging), as demonstrated in Table 6. By introducing our auxiliary neural network, our NuSegDG framework achieves remarkable generalization-efficiency trade-offs.
In addition to the ablation study on the individual components of NuSegDG, we also conduct a supplementary experiment to compare the traditional point prompt approach used in the original SAM with our proposed density map prompt. As shown in Table 7, the results indicate that while increasing the number of point prompts from one to seven can lead to slight improvements in segmentation performance, the use of the density map prompt consistently achieves the best results across all metrics. For instance, the density map prompt yields a Dice score of 93.17%, an mIoU of 87.46%, and an F1 score of 93.35%, while substantially reducing the Hausdorff Distance (HD) to 33.18, which is significantly lower than those obtained by any point prompt configuration. This experiment demonstrates that the density map prompt not only captures richer semantic and positional cues compared to single or multiple point prompts but also substantially improves segmentation accuracy and boundary adherence. Consequently, this reinforces the effectiveness of our Gaussian-Kernel Prompt Encoder in providing robust and efficient guidance for segmentation decoding, so that it contributes to the overall domain generalization capability of the NuSegDG framework.
4.6 Analysis of Hyper-Parameters
In this section, we perform a comprehensive hyper-parameters analysis of our NuSegDG model. As reported in Section 3.2 and 3.3, NuSegDG contains two hyper-parameters, including the Gaussian kernel size $r$ in GKP-Encoder and the number of heterogeneous space $N$ in HS-Adapter. For the kernel size, we perform a grid search under the fully-supervised learning to select an optimal configuration. Fig. 4a shows the average Dice and mIoU of NuSegDG on the four nuclei domains with different kernel sizes. It is indicated that the NuSegDG with $r=10$ demonstrates the best performance due to the sufficient semantic information prompts. However, excessive kernel size may generate false positive errors, which cannot offer additional benefits. For the number of heterogeneous space, we provide the result of grid search in Fig. 4b. We observe that the NuSegDG with $N=2$ obtains the best performance. Setting more heterogeneous space significantly increases the computational complexity of NuSegDG, which is not suitable for limited training samples in nuclei domains. These experimental results prove the importance of tuning these hyper-parameters to improve the efficiency of our NuSegDG framework in learning domain-specific knowledge. In addition, compared to existing fully fine-tuned methods (e.g., MedSAM [ma2024segment], SAMMI [huang2024segment]), NuSegDG only requires a small number of trainable parameters, significantly reducing computational costs. Moreover, the standard SAM [kirillov2023segment] and other variants (e.g., LeSAM [gu2024lesam], SAM-CL [zhong2024convolution], Med-SA [wu2023medical]) need multiple point prompts to generate accurate segmentation results. On the contrary, NuSegDG can realize better performance with a single-point prompt, which is more effortless and helps reducing annotation time and fatigue. Further, we fine-tune the Auxiliary Network (AN) on four nuclei instance segmentation datasets. The result has been presented in Table 8. Our findings indicate that fine-tuning AN yields a modest improvement in performance across several metrics (e.g., AJI, DQ, SQ, and PQ) compared to using the pretrained weights. This suggests that while fine-tuning can help further optimize the performance, the pretrained model already captures critical features required for generating accurate instance maps. Consequently, our initial design choice of using pretrained weights remains justified in terms of computational efficiency, with the fine-tuned version providing an upper-bound performance reference.